nini01:

3(−4x) + 3(5) = 12 −12x + 15 = 12 −12x + 15 − 15 = 12 − 15 −12x = −3 x =

TYRANNNNNNNNN:

What are we trying to answer here

nini01:

x

darkknight:

divide both sides by -12

n32carrillo:

x=4 because if you multiply 4 by -3 you get 12. also, you divide -3 from both sides and u still get 4.

boredfr:

3(−4x)+(3)(5)=12 Step 1: Simplify both sides of the equation. −12x+15=12 Step 2: Subtract 15 from both sides. −12x+15−15=12−15 −12x=−3 Step 3: Divide both sides by -12. −12x −12 = −3 −12 x= 1 4

boredfr:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @boredfr 3(−4x)+(3)(5)=12 Step 1: Simplify both sides of the equation. −12x+15=12 Step 2: Subtract 15 from both sides. −12x+15−15=12−15 −12x=−3 Step 3: Divide both sides by -12. −12x −12 = −3 −12 x= 1 4 \(\color{#0cbb34}{\text{End of Quote}}\) 1/4